the altitude of a triangle is 4m longer than the base. If its area is 126 square meter, find the length of the altitude.
can somebody answer this problem..thank you!

Answer by mananth(16946)

base =x
altitude= (x+4)
1/2*x*(x+4)=126
x^2+4x=252
x^2+4x-252=0
x^2+18x-14x-252=0
x(x+18)-14(x+18)=0
(x+18)(x-14)=0
x= 14m which is positive.
base = 14 height = 14+4=18m